The length of a rectangle is 5 meters longer than the width.Find the length and width if the perimeter is 54 meters.
Answer by nadinadan(32)

So let's say that the length is L and the width is W, of the rectangle.
The problem says that L=5+W, the length is 5 more than width.
We know that the perimeter of the rectangle is 2L+2W,and the problem says that the perimeter is 54, so 2L+2W=54.
We have 2 equations now,
L=5+W and
2L+2W=54 if we replace the first equation into the second one we have
2(5+W)+2W=54, that makes 10+2W+2W=54 If we simplify, 4W=54-10 so 4W=44 amd W=44/4 which makes W=11.
Now we replace W in the first equation L=5+W, so L=5+11 so L=16
To verify ourselves, we replace L and W in the perimeter equation.
2L+2W=54 2*16+2*11=54 32+22=54, so we did good.
Let me know if you need more help.